### v1

# MCCoder: Streamlining Motion Control with LLM-Assisted Code Generation and Rigorous Verification

Abstract
Large Language Models (LLMs) have shown considerable promise in code generation. However, the automation sector, especially in motion control, continues to rely heavily on manual programming due to the complexity of tasks and critical safety considerations. In this domain, incorrect code execution can pose risks to both machinery and personnel, necessitating specialized expertise. To address these challenges, we introduce MCCoder, an LLM-powered system designed to generate code that addresses complex motion control tasks, with integrated soft-motion data verification. MCCoder enhances code generation through multitask decomposition, hybrid retrieval-augmented generation (RAG), and self-correction with a private motion library. Moreover, it supports data verification by logging detailed trajectory data and providing simulations and plots, allowing users to assess the accuracy of the generated code and bolstering confidence in LLM-based programming. To ensure robust validation, we propose MCEVAL, an evaluation dataset with metrics tailored to motion control tasks of varying difficulties. Experiments indicate that MCCoder improves performance by 11.61% overall and by 66.12% on complex tasks in MCEVAL dataset compared with base models with naive RAG. This system and dataset aim to facilitate the application of code generation in automation settings with strict safety requirements. MCCoder is publicly available at https://github.com/MCCodeAI/MCCoder.
I INTRODUCTION
Figure 1: Naive RAG code generation vs MCCoder. Code generation with Naive RAG has a low pass rate, requiring extensive debugging and posing safety concerns. In contrast, MCCoder generates code through task decomposition, correction, and simulation with soft-motion, ensuring data verification and safe machine deployment.
Motion control, a cornerstone of automation, has significantly advanced industrial processes since the advent of NC machines in the 1950s. By precisely controlling actuators, it has revolutionized manufacturing, from early CNC machines to today’s sophisticated AI-driven robots and semiconductor equipment. While motion control encompasses various programming methods (e.g., CAD/CAM, PLC, robot teaching), this paper focuses on the complex domain of motion API invocation. Today, software engineers in this industry still manually program due to its complexity and safety concerns. Some private motion libraries have over 1,000 APIs with various arguments, creating a steep learning curve and making reliance on automatic programming tools difficult. Additionally, engineers must code and iteratively test with electrical and mechanical components to ensure optimal motion performance and eliminate potential hazards for humans and machines during execution.
The recent spotlight on LLM-based code generation has unveiled impressive ability to address programming challenges [1]. Although traditional code generation in automation control often relied on model-driven or rule-based approaches, recent advancements have started to explore LLM-based methods. For example, a retrieval-augmented approach [2] was proposed to generate IEC 61131-3 programs, while LLM4PLC [3] used user feedback and external verification tools to guide the generation of LLM-based PLC codes.
However, current attempts to use LLMs for motion control primarily focus on standardized PLC languages, neglecting higher-level languages like Python or C++ commonly used in complex automation equipment. Furthermore, there is a lack of thorough research on verifying the safety and effectiveness of generated control code, as well as a shortage of suitable datasets for evaluation.
To address these limitations, we propose the MCCoder, which utilizes LLMs to generate Python code for motion control based on natural language instructions. This system employs a private motion library to decompose complex tasks, generate code, and use soft-motion for simulation running and data logging for verification. In addition, we build the MCEVAL dataset to evaluate the generated code. Figure 1 highlights a comparation between LLM-based code generation with naive RAG and LLM-powered MCCoder. Experiments demonstrated that MCCoder enhances code generation performance, achieving an overall improvement of 11.61% and a significant increase of 66.12% in complex tasks in the MCEVAL dataset.
The primary contributions of our work are as follows.
  • Code Generation System: We introduce MCCoder system that effectively generates Python code for motion control using LLMs. It incorporates task decomposition, retrieval, code generation, soft-motion, self-correction and data verification to ensure code safety and effectiveness in addressing complex real-world programming challenges. Soft-motion is carefully selected as a flexible motion controller integrated with LLMs for rapid simulation, execution, self-correction, and rigorous verification.
  • Evaluation Dataset: Current code generation datasets and metrics are not designed for the motion control domain. We present MCEVAL, the first benchmark dataset and metrics specifically tailored for evaluating LLM-powered motion control code. This resource fills a gap in the field.
  • Benchmarking: Through extensive experiments comparing state-of-the-art LLMs with MCCoder, we offer insights into their performance and evaluate their suitability for deployment in industrial motion control applications.
II PRELIMINARY
MCCoder utilizes a soft-motion architecture to generate Python code for motion control by invoking APIs from a private library.
II-A Soft-motion
The soft-motion controller has a flexible and lightweight architecture that runs on a general PC without the need for specialized hardware. It uses a dedicated CPU core with a real-time OS and software motion algorithms to deliver performance comparable to high-end hardware controllers. MCCoder sends control code to the soft-motion engine through a DLL. Soft-motion executes the code in a simulation engine first, allowing for data logging and verification before deployment. Additionally, its PC-based software architecture makes soft-motion ideal for integrating new functions, including LLMs and other AI technologies.
II-B Control Code and Private Library
Unlike standardized PLC programming like IEC 61131-3, control codes programming in Python are more flexible and difficult to apply a fixed template. In this research, we use a private motion library in the soft-motion controller with more than 1,000 APIs capable of controlling up to 128 axes and 256 independent task channels. These APIs encompass most motion control functionalities, such as fieldbus communication, digital and analog IO control, and various trajectory and position/velocity/current control for servo axes. The private library also includes hundreds of sample codes. Section II-B shows a simple code example to move Axes to positions, including the INITIALIZE, main control code and CLOSE procedure.
Example 1. A Motion Control Code Demo   Inputs:   $Axes$ = [$A_{i}$] // Axes objects from LLMs   $IO_{Ins}$ = [$I_{i}$]; // Input objects from LLMs   $IO_{Outs}$ = [$O_{i}$] // Output objects from LLMs   procedure INITIALIZE():       $MCLib.CreateDevice()$       $MCLib.StartCommunication()$       $MCLib.SetandStartLog(Axes,IO_{Ins},IO_{Outs})$   // Main control code   for $i$ in $Axes$ do       $posCommand=MotionPosCommand()$      $posCommand.profile.type=ProfileType_{i}$      $posCommand.axis=i$      $posCommand.target=Pos_{i}$      $MCLib.motion.StartPos(posCommand)$      $MCLib.motion.Wait(i)$   end for   procedure CLOSE():           $MCLib.CloseDevice()$   end procedure   Return: Success or error code
III METHODOLOGY
We developed the MCCoder system to tackle the code generation challenges in motion control. Figure 2 shows an overview of the MCCoder system. It involves six modules: task decomposition, retrieval, code generation, soft-motion, self-correction and data verification.
Figure 2: Overview of the MCCoder system. It integrates multitask decomposition, hybrid retrieval, code generation, and self-correction with a private motion library to enhance performance, offering motion data visualization for execution verification before deployment. Soft-motion is a key component in the system to streamline motion control with LLM-assisted code generation with rapid execution and self-correction cycle as a rigorous verification, which is critical for control code programming.
III-A Task Decomposition
The MCCoder system receives user questions as a motion task. It could be a single task or a multitask composing with many sub-tasks. This module will read the task and decompose them to be a list of sub-tasks.
III-B Retrieval
The retrieval module extracts APIs and expert-created sample codes from the private library, each accompanied by clear instructions. Retrieval methods are commonly employed in code generation to provide relevant samples and contexts. Various Retrieval-Augmented Generation (RAG) techniques are available, including native RAG, advanced RAG, and modular RAG [4]. Given that the library contains uniformly formatted sample codes and APIs, we employ Sparse Retrieval (BM25), Dense Retrieval (embedding-based VectorStores), and a re-ranker to match user queries with the most relevant content, setting the top-$k$ as 10.
III-C Code Generation
The code generation model processes sub-tasks and retrieved sample codes to generate sub-codes based on prompts, repeating this for m tasks. It then combines the user question and sub-codes to generate the final code. A critical but often overlooked aspect is prompt engineering. To consistently refine the prompts, we iterated and established 8 guidelines to help models avoid common mistakes.
III-D Soft-Motion
Soft-motion includes a real-time engine for operating actual machines and a simulation engine. MCCoder sends the final code to the soft-motion controller, which runs the code in the simulation engine and returns execution feedback.
III-E Self-Correction
If the soft-motion engine returns a syntax or API error, MCCoder uses the error to retrieve relevant documents, regenerate code and run it again for self-correction iteratively. If no error is detected, soft-motion logs the entire trajectory data for subsequent verification.
III-F Data Verification
We can design a 3D model of a machine and import the model file into the 3D simulator, which displays the motion in real-time, providing users with an intuitive verification when it runs. Additionally, logging data is recorded every 1ms by the soft-motion engine as the process data. These features are essential to build user trust in an AI-driven system through visual and data-driven verification.
IV EVALUATION DATASET
Existing code generation datasets, like HumanEval [5] and MBPP EvalPlus [6], focus on general-purpose code evaluated via unit tests. The generation of control code requires execution in motion controllers to record the endpoint and trajectory data. MCEVAL evaluation dataset fills this gap for the automation industry. We meticulously constructed MCEVAL, verifying each task both manually and through soft-motion simulations to ensure that the programming tasks are well-formed.
IV-A Construction
MCEVAL comprises 116 tasks that were chosen to provide a diverse but manageable set of motion control programming challenges. It offers substantial diversity, covering most common used motion control functions in the library. These functions include point-to-point motion, linear, circular, and helical interpolation, splines, short-segment look-ahead, compensation, and event-driven interactions of axes and IOs under various profiles, as well as their combinations. Each task includes a TaskId, Instruction, CanonicalCode, and Difficulty. The properties of these tasks are detailed in Table I.
To mimic a human engineer’s problem-solving capabilities, we structured the tasks into three difficulty levels. These levels are differentiated by sample code, primarily demonstrating single API function calls and not covering all argument combinations (e.g., there are 19 acceleration profiles). The examples are shown in Sections IV-A and IV-A. Difficulty Level 1 involves invoking similar APIs as in the sample codes but with different numerical arguments, which should be easy for most LLMs if the retrieved sample codes are correct. Difficulty Level 2 uses arguments not shown in the sample codes, making it more challenging. LLMs might make mistakes but should utilize error information or retrieved content for self-correction. Difficulty Level 3 involves combining up to 8 different motion tasks. Consolidating more than 8 tasks into a single instruction may exceed input or output token limits for most models. These levels emulate real-world scenarios from easy to complex motion tasks.
TABLE I: Property and Structure of MCEVAL Dataset
Tasks Numbers Instruction Length Cannonical Code Length
Difficulty Level 1 61 267 2268
Difficulty Level 2 32 289 1351
Difficulty Level 3 23 406 2926
Total 116 273 2145
Example 2. Sample Code   Instruction: Move Axis 2 to 100, velocity as 1000.   Return:       $posCmd=MotionPosCommand()$       $posCmd.profile.type=Trapezoidal$       $posCmd.axis=2$       $posCmd.target=100$       $posCmd.profile.velocity=1000$       $MCLib.motion.StartPos(posCmd)$
[…]
IV-B Metrics
To comprehensively assess the generation of code on the evaluation dataset, the following metrics are chosen based on control expertise.
IV-B1 First Time Pass Rate (FTPR)
FTPR is calculated as Equation 1 representing the proportion of codes that pass the test on the first attempt [7]. Pass@k is commonly used in code generation metrics and repeats many times to eliminate bias, but practical users prefer code that works correctly on the first try. We chose this straightforward measure to align with real-world use cases for control code.
$FTPR=\frac{N_{Passed}}{N_{Total}}*100$ (1)
IV-B2 MatchEndPoints and DTW
The soft-motion engine generates log data every 1ms during code execution. Log files from the canonical code in MCEVAL dataset and the generated code are compared. MatchEndPoints compares only the end points, while Dynamic Time Warping (DTW) compares all trajectory points, measuring the similarity of time series data. In motion control, it depends on specific scenario whether to check end points or trajectories, for instance, interpolation motions should match trajectories, while point-to-point motions only need endpoint matching. We will measure both FTPR (MatchEndPoints) and FTPR (DTW).
V EXPERIMENTS
V-A Experimental Settings
We access closed-source models through APIs and open-source one on a local server with eight NVIDIA GeForce RTX 3090 GPUs and Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz. A soft-motion controller and 3D simulator were installed on a Windows 11 PC, ready for code execution and simulation.
V-B Base Models
LLMs evolve rapidly, continuously setting new state-of-the-art benchmarks. For this study, we selected the latest closed-source and open-source LLMs as base models. Given the average token length of 1,424 in the sample codes, we opted for models with relatively large context windows and output lengths, as detailed in Table II. GPT-4o, GPT-4o-mini, Lama-3.1-70b, and DeepSeek-Coder-v2 were selected for their proven performance in code generation benchmarks.
TABLE II: Base Models for experiments
Model Context Window Max Output Tokens Active Params
gpt-4o 128,000 4,096 -
gpt-4o-mini 128,000 16,384 -
llama-3.1-70b 128,000 2,048 70B
deepseek-coder-v2 128,000 4,096 21B
V-C Baseline
Since the private library is not in pre-trained LLM data, zero-shot evaluation is inapplicable. We established a baseline using the base model with naive RAG and top-$k$ = 10, comparing it to the base models integrated into the MCCoder system.
V-D Experimental Results
TABLE III: Experiments results of various models in MCEVAL datasets. For each model, the baseline naive RAG and the MCCoder system were compared. Metrics include FTPR (MatchEndPoints) and FTPR (DTW), with the latter being more stringent. The overall first-time pass rate was assessed and performance was evaluated at three difficulty levels. L1, L2, and L3. The best performance are boldfaced for Naive RAG, and underlined for MCCoder.
Model Method FTPR(MatchEndPoints) $\uparrow$ FTPR(DTW) $\uparrow$
OVERRALL L1 L2 L3 OVERRALL L1 L2 L3
gpt-4o Naive RAG 67.24 78.69 59.38 47.83 53.45 63.93 46.88 34.78
MCCoder 75.86 83.61 65.62 69.57 67.24 73.77 59.38 60.87
gpt-4o-mini Naive RAG 60.34 75.41 46.88 39.13 43.97 52.46 37.50 30.43
MCCoder 66.38 75.41 53.12 60.87 49.14 55.74 37.50 47.83
llama-3.1-70b Naive RAG 59.48 73.77 50.00 34.78 48.28 59.02 40.62 30.43
MCCoder 64.66 68.85 59.38 60.87 50.00 52.46 50.00 43.48
deepseek-coder-v2 Naive RAG 67.24 77.05 62.50 47.83 56.90 68.85 50.00 34.78
MCCoder 75.86 80.33 65.62 78.26 61.21 67.21 53.12 56.52
We compared naive RAG and MCCoder using FTPR (MatchEndPoints) and FTPR (DTW) metrics. The overall pass rates and three difficulty levels were also evaluated for each model, as shown in Table III. We will present results with research questions.
V-D1 RQ1. Does MCCoder imporve the code generation performance?
MCCoder yielded an overall FTPR improvement of 11.61%. While this might appear modest at first glance, a deeper analysis reveals a more substantial impact at higher difficulty levels. At Level 3, the improvement increases to 66. 12%, demonstrating MCCoder’s exceptional ability to tackle complex real-world challenges.
V-D2 RQ2. Which base model behaves the best with MCCoder?
GPT-4o demonstrated the best overall performance, particularly with MCCoder. However, DeepSeek-Coder-v2 exhibited very strong performance comparable to GPT-4o and even outperformed it in naive RAG, which is surprising given its 24 billion active parameters. In the HumanEval evaluation, DeepSeek-Coder-v2 scored 90.2, close to GPT-4o’s 91.0 [8]. Our experiments show similar results. GPT-4o-mini and Lama-3.1-70B performed slightly weaker.
V-D3 RQ3. Why is MCCoder’s data verification critical for motion control?
In general-purpose Python programming, verification typically focuses on execution results via unit tests. In motion control, however, engineers are concerned with the entire execution process, including the trajectory and endpoints of the objects under control to confirm if it is correct and safe. LLMs’ inferencing and code generation capabilities are often perceived as a black-box, raising transparency concerns. Therefore, traceable data that clearly represent the entire process is essential. For each task execution, MCCoder generates a logging file, capturing the entire process as time-series data for evaluation. In real-world scenarios, 1D-3D plots of Axes and IOs, along with real-time 3D simulator of the machine model, provide a multimodal, traceable method for verifying whether the trajectory and endpoints align with expectations. MCCoder provides traceability that is crucial for building trust and enabling deployment, as presented in Figure 3.
Figure 3: Example of logging data and plots from MCCoder system, providing data and visual virification for users.
V-E Error Analysis
During the evaluation of the control codes, the soft-motion engine returns some errors. We analyze the errors in code generation with gpt-4o and their causes in detail.
  1. 1.
    API errors (5.2%): These errors indicate incorrect API function invoking. But the API error rate is relatively small, showing the effectiveness of the current retrieval method for relevant sample codes reference.
  2. 2.
    Argument errors (79%): The most prevalent error type involves correct API invocations with incorrect arguments, frequently stemming from hallucinations or disregard for prompt instructions. Self-correction can sometimes address and correct these errors. Explicit error messages like ”type object ProfileType has no attribute scurve. Did you mean: SCurve?” significantly enhance the likelihood of successful correction in subsequent attempts. In contrast, implicit error messages hinder accurate identification and correction.
  3. 3.
    Syntax errors (15.8%): These errors predominantly arise from hallucinations or the invention of nonexistent functions and libraries, resulting in syntactic inaccuracies.
V-F Ablation Study
We conducted an ablation study focusing on prompt engineering, task decomposition, and self-correction modules with the base models GPT-4o and GPT-4o-mini. As shown in Figure 4, prompt engineering plays a significant role in the MCCoder system, causing a 20.51% performance drop in FTPR (DTW) for GPT-4o and a 15.79% drop for GPT-4o-mini, with a larger impact on the larger model. Task decomposition resulted in a performance drop of around 13% for both models, indicating that more than half of the difficulty level 3 tasks failed without this module. The self-correction module had a relatively smaller impact, with a 1-5% performance difference. However, as we improve the precision of the soft-motion controller error information in the future, this module is expected to play a more substantial role.
Figure 4: Ablation study of MCCoder. Prompt engineering and task decomposition play critical roles in the system, while self-correction module shows a relatively small impact because of accuracy of the error information.
VI DISCUSSION
VI-A Engineers’ feedback
We presented the system to a group of control engineers, ranging from novices to experts, and collected their feedback. Both groups highly regarded MCCoder for facilitating learning new functions and improving coding efficiency. They highlighted the necessity of human-involved verification to address safety concerns, an issue they encountered when using general LLMs for control code generation. Moreover, they suggested enhancing code interpretability, such as through visual flowcharts or explanations of API usage, to assist in better understanding and debugging when minor errors occur. This feedback has provided valuable insights for future improvements and deployment.
VI-B Limitations
The current MCEVAL dataset may not comprehensively encompass the full spectrum of motion control tasks, necessitating further testing across a broader range of motion scenarios to enhance the dataset. Additionally, the error messages returned by the soft-motion controller are occasionally ambiguous, which impedes MCCoder’s self-correction capabilities. While the system’s effectiveness has been demonstrated through the four existing model evaluations, conducting additional model tests could yield further insights for refinement and improvement.
VII RELATED WORK
VII-A Code Generation with LLMs and Strategies
Recent advancements in LLMs, such as OpenAI’s Codex, GitHub’s Copilot, Google’s Gemini, and Meta’s CodeLlama, have significantly enhanced code generation, improving both productivity and code quality [9]. Trained on diverse codebases with advanced NLP techniques, these models excel in code completion, refinement, and debugging [10]. Benchmarks like HumanEval, MBPP, and CodeXGLUE reveal that these models are nearing or exceeding human-level performance in coding tasks [11]. Enhanced model architectures with larger context windows and sizes allow them to tackle more complex tasks [12]. Code generation strategies in LLMs include task planning [13], in-context learning [14], Chain-of-Thought prompting [15], RAG, and post-process methods like self-consistency, DIN-SQL [16], and DAIL-SQL [17], all of which improve the efficiency, accuracy, and reliability of these models.
VII-B Code Generation in Motion Control
Experts in industrial automation have explored methods to automate programming tasks, including model-driven development environments [18] and rule-based systems like automated Matlab-to-C++ translators [19]. 13 code generation methods for control logic has been classified since 2004 [20], but these traditional methods were limited to small-scale applications. Large Language Models (LLMs) offer significant advantages for large-scale applications. Researchers are now using LLMs for innovative code generation methods. For example, a retrieval-augmented approach was proposed for generating IEC 61131-3 Structured Text programs using GPT-4, LangChain, and OpenPLC, validated through expert simulations [2]. Similarly, LLM4PLC was developed as a user-guided iterative pipeline using syntax checkers and LLM fine-tuning for PLC programming [3]. These methods represent a promising start for automating control code generation.
VIII CONCLUSION
In this paper, we introduce MCCoder to leverage LLMs for generating motion control code. MCCoder addresses the complexities and safety-critical aspects of motion control programming by integrating prompt engineering, task decomposition, retrieval, code generation, soft-motion and verification through simulation and data logging. Validated using the MCEVAL dataset, MCCoder significantly improves code generation performance, especially for the most challenging tasks. These advancement provide valuable insights and guidelines for future deployment and research in industrial automation and control code generation. Future work will focus on expanding the MCEVAL dataset, improving error message clarity from the soft-motion controller, and exploring fine-tuning with larger models to further enhance MCCoder’s capabilities.
ACKNOWLEDGMENT
This work is funded by National Natural Science Foundation of China Grant No. 72371271, the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), the Nansha Key Area Science and Technology Project (No. 2023ZD003), and Project No. 2021JC02X191.